When 10g of a mixture of potassium chloride and anhydrous sodium sulphate is dissolved in water and excess barium chloride solution added, 6.9g of barium sulphate is precipitated. Calculate the composition of the mixture
1
Expert's answer
2020-08-25T12:14:37-0400

Upon addition of the barium chloride to a mixture of potassium chloride and sodium sulphate, only sodium sulphate reacts with barium chloride. The balanced equation of the reaction between sodium sulphate and barium chloride is:

Na2SO4 + BaCl2 "\rightarrow" 2NaCl + BaSO4"\downarrow" .

According to the stoichiometric coefficients, the number of the moles of sodium sulphate that react and the number of the moles of barium sulphate that is formed relate as:

"n(Na_2SO_4) = n(BaSO_4)".

The number of the moles of barium sulphate can be calculated from its mass using its molar weight "M(BaSO_4) = 233.39" g/mol:

"n(BaSO_4) = \frac{m}{M} = \frac{6.9 \text{ g}}{233.39 \text{ g/mol}} =0.02956" mol.

Therefore, the number of the moles of sodium sulphate and its mass ("M(Na_2SO_4) = 142.04" g/mol) are:

"n(Na_2SO_4) = 0.02956" mol,

"m(Na_2SO_4) = n\u00b7M"

"m(Na_2SO_4)= 0.02956\text{ mol}·142.04\text{ g/mol} = 4.2" g.

Therefore, the composition of the mixture is the following: 4.2 g of Na2SO4 and 10-4.2 = 5.8 g of KCl.

Answer: 4.2 g of Na2SO4 and 5.8 g of KCl.